20.16 - Current A 14.0-gauge copper wire has a diameter of 1.628 mm. What magnitude current flows when the drift velocity is 1.41 mm/s? (See Example 20.3 for useful information.) 22.2A Submit Anser Incorrect. Tries 3/10 Previous Tries
Answer #1

drift velocity = I/nAq

where I = current

n = number of atoms

A = area

q =charge on electron

0.00141 = I / (14/63.5 * 6.023*10^23) * (pi/4 * 0.001628^2) * 1.6*10^-19

I = 6.24*10^-5 A

= 62.4 uA